OpenStudy (anonymous):

Solve (1-4xy^2)(dy/dx)=y^3 by regarding y as the independent variable rather than x?????

OpenStudy (anonymous):

\[\begin{align*} (1-4xy^2)\frac{dy}{dx}&=y^3\\\\ \frac{dy}{dx}&=\frac{y^3}{1-4xy^2}\\\\ \frac{dx}{dy}&=\frac{1-4xy^2}{y^3}\\\\ \frac{dx}{dy}&=\frac{1}{y^3}-\frac{4x}{y} \end{align*}\] Now you have an ODE linear in \(x\).